OpenStudy (anonymous):

use implicit differentiation to find dy/dx: y^2-2xy=0

OpenStudy (tkhunny):

What's preventing you from diving in and completing this task? Please demonstrate your initial efforts.

OpenStudy (anonymous):

add 2xy to both sides then take derivate of both sides (y^2)^1=(2xy)^1

OpenStudy (tkhunny):

You have not ever seen such an example. We are assuming that y = f(x). We have some Chain Rules in our future. \(Given\;y = f(x),\;we\;have\;\dfrac{d}{dx}y^{2} = 2y\dfrac{dy}{dx}\) Okay, you do the piece that requires the Product Rule and the Chain Rule. \(Given\;y = f(x),\;we\;have\;\dfrac{d}{dx}(-2xy) = -2\dfrac{d}{dx}(xy) =\;??\)

OpenStudy (tkhunny):

Please don't do what @BlackLabel just suggested. That would be quite incorrect.